# Polynormer: Polynomial-Expressive Graph Transformer in Linear Time

### 02 Mar 2024 v1

@@ 3.1 A Polynomial-Expressive Base Model with Attention
As any polynomial can be expressed as a linear combination of monomials, a natural way to build a polynomial-expressive model is to explicitly capture the monomial terms. To this end, we follow conventional polynomial networks (Chrysos et al., 2020; 2022; Wang et al., 2021) to provide Definition 3.1 of the polynomial expressivity.
@@ 4 Experiments
We have conducted an extensive evaluation of Polynormer against state-of-the-art GNN and GT models on both homophilic and heterophilic graphs, where nearby nodes tend to have the same or different labels. Besides, we demonstrate the scalability of Polynormer on large-scale graphs that contain millions of nodes. In addition, we perform additional ablation analysis to understand the effectiveness of the local-to-global attention scheme adopted by Polynormer. Finally, we visualize Polynormer attention scores to showcase its ability of capturing critical global structures.
Experimental Setup. Appendix E shows the details of all $13$ datasets used in our experiments, which consist of $7$ homophilic and $6$ heterophilic graphs. Notably, the heterophilic datasets are from Platonov et al. (2023); Lim et al. (2021), which have addressed serious issues (e.g., train-test data leakage) of conventional popular datasets such as chameleon and squirrel. We compare Polynormer against $16$ GNNs that include state-of-the-art models on our chosen datasets. Note that prior graph polynomial models except tGNN (Hua et al., 2022) discussed in Section 2 run out of GPU memory even on the smallest dataset in our experiments. Besides, we also evaluate $6$ GTs that have shown promising results on the node classification task. We report the performance results of baselines from their original papers or official leaderboards whenever possible, as those results are obtained by well-tuned models. For baselines whose results are not publicly available on given datasets, we tune their hyperparameters to achieve the highest possible accuracy. Detailed hyperparameter settings of baselines and Polynormer are available in Appendix H. Our hardware information is provided in Appendix F.
@@ 4.1 Performance on Homophilic and Heterophilic Graphs
Performance on homophilic graphs. We first compare Polynormer with $16$ popular baselines that include state-of-the-art GNNs or GTs on $5$ common homophilic datasets. As shown in Table 1, Polynormer is able to outperform all baselines on $3$ out of $5$ datasets, which clearly demonstrate the efficacy of the proposed polynomial-expressive model. Moreover, incorporating the $ReLU$ function can lead to further improvements in the accuracy of Polynormer. This is because the nonlinearity imposed by $ReLU$ introduces additional higher-order monomials, which in turn enhance the quality of node representations. It is noteworthy that the accuracy gain of Polynormer over state-of-the-art baselines is $1.65\%$ on Computer, which is a nontrivial improvement given that those baselines have been finely tuned on these well-established homophilic datasets.
Performance on heterophilic graphs. Table 2 reports the average results over $10$ runs on heterophilic graphs. Notably, the GT baselines underperform state-of-the-art GNNs on all $5$ datasets, which raises concerns about whether those prior GT models properly utilize the expressivity brought by the self-attention module. Moreover, there is no baseline model that consistently ranks among the top $3$ models across $5$ datasets. In contrast, by integrating the attention mechanism into the polynomial-expressive model, Polynormer surpasses all baselines on $4$ out of $5$ datasets. Furthermore, Polynormer-r consistently outperforms state-of-the-art baselines across all datasets with the accuracy improvement by a margin up to $3.55\%$ (i.e., $97.46\%$ vs. $93.91\%$ on minesweeper).
@@ 4.2 Performance on Large Graphs
We further conduct experiments on large graphs by comparing Polynormer against scalable GTs as well as some representative GNN models. Since the graphs in ogbn-products and pokec are too large to be loaded into the GPU memory for full-batch training, we leverage the random partitioning method adopted by prior GT models Wu et al. (2022; 2023) to perform mini-batch training. As shown in Table 3, Polynormer-r outperforms all baselines on these large graphs with the accuracy gain of up to $4.06\%$. Besides, we can observe that the accuracy of Polynormer drops $1.64\%$ when removing $ReLU$ activation on ogbn-arxiv. Thus, we believe the non-linearity associated with $ReLU$ is more critical on ogbn-arxiv, which is known to be a challenging dataset (Shirzad et al., 2023). In addition, Appendix J provides the training time and memory usage of Polynormer and GT baselines. Furthermore, we assess Polynormer using an industrial-level graph benchmark named TpuGraphs (Phothilimthana et al., 2024), and the details are provided in Appendix L.
@@ 4.3 Ablation Analysis on Polynormer Attention Schemes
Figure 3 shows the comparison of the state-of-the-art (SOTA) baseline (red bar), Polynormer without global attention (orange bar), a variant of Polynormer where the local and global attention modules are applied simultaneously to update node features per layer (blue bar), and Polynormer (green bar). We provide the detailed architecture of the Polynormer variant in Appendix G. Besides, we omit the results of Polynormer without local attention, as it completely ignores the graph structural information and thus performs poorly on the graph datasets.
@@ Appendix C Polynomial Expressivity of Prior Graph Models
Conclusion. Due to the limitations of polynomial expressivity analyzed above, previous GTs and high-order GNNs still necessitate non-linear activation functions to attain reasonable accuracy. In contrast, our method achieves higher accuracy than these models even without activation functions, thanks to its high polynomial expressivity.
@@ H.2 Polynormer
Warm-up Epochs Total Epochs Local Layers Global Layers Dropout
@@
Photo $200$ $2000$ $7$ $2$ $0.7$
CS $100$ $500$ $5$ $2$ $0.3$
Physics $100$ $500$ $5$ $2$ $0.5$
WikiCS $100$ $500$ $7$ $2$ $0.5$
@@
questions $200$ $1500$ $5$ $3$ $0.3$
ogbn-arxiv $500$ $2000$ $7$ $2$ $0.5$
ogbn-products $100$ $500$ $10$ $2$ $0.5$
pokec $500$ $1000$ $7$ $2$ $0.3$
Like the baseline models, we set the hidden dimension to $512$ and the learning rate to $0.001$. Additionally, we introduce a warm-up stage dedicated to training the local module. This step ensures that the node representations generated by the local module capture meaningful graph structural information before being passed to the global module. Moreover, we leverage $8$ heads for our attention modules. For the ogbn-products and pokec datasets, we use minibatch training with a batch size of $250,000$ and $600,000$ respectively, while for the other datasets, we employ full batch training. In Table 5, we provide the detailed hyperparameters for Polynormer used with each dataset.
@@ Appendix I Model Implementation
To validate the linear complexity of Polynormer with respect to the number of nodes/edges in a graph, we generate synthetic Erdos-Renyi (ER) graphs with node count ranging from $1,000$ to $200,000$. We control the edge probability in the ER model to achieve an average node degree of around $5$ and set the node feature dimension to $100$. Figure 5 (left) clearly shows that the memory usage of Polynormer linearly increases with graph size, confirming its linear space complexity. As for the time complexity of Polynormer, we observe in Figure 5 (right) that the GPU training time remains nearly constant when increasing the graph size. We attribute it to that Polynormer is essentially a compute-bound model, and the GPU device (RTX A6000) used in our experiments can efficiently handle parallel computations on graphs with different sizes. This observation is further supported by the CPU training time results, which linearly increase with graph size, as the CPU is less powerful than the GPU for parallel computation.
@@
ogbn-arxiv ogbn-products pokec Method Train/Epoch (s) Mem. (GB) Train/Epoch (s) Mem. (GB) Train/Epoch (s) Mem. (GB) GAT 0.16 10.64 0.97 9.04 1.36 9.57 GraphGPS $\underline{1.32}$ $\underline{38.91}$ $OOM$ $OOM$ $OOM$ $OOM$ NAGphormer $4.26$ 5.15 $9.64$ 7.91 $38.32$ 6.12 Exphormer 0.74 $\underline{34.04}$ $OOM$ $OOM$ $OOM$ $OOM$ NodeFormer $\underline{1.20}$ $\underline{16.30}$ $3.37$ $31.55$ $6.15$ $17.21$ DIFFormer $\underline{0.77}$ $\underline{24.51}$ 1.50 $16.24$ 3.93 $16.03$ GOAT $12.32$ 6.98 $20.11$ 9.64 $51.93$ 8.73 Polynormer 0.31 $\underline{16.09}$ 3.13 $12.91$ 2.64 $11.54$
@@
Apart from comparing Polynormer against conventional GRL models on mainstream graph datasets, we further evaluate it on TpuGraphs (Phothilimthana et al., 2024), a large-scale runtime prediction dataset on tensor computation graphs.

### 06 Apr 2024 v3

@@ 3.1 A Polynomial-Expressive Base Model with Attention
By following the concept of polynomial networks in Hua et al. (2022); Chrysos et al. (2020; 2022), we provide Definition 3.1 of the polynomial expressivity, which measures the capability of learning high-degree polynomial functions that map input node features into output node representations. Appendix M provides a detailed discussion on our definition and comparison to prior work.
@@ 4 Experiments
We have conducted an extensive evaluation of Polynormer against state-of-the-art (SOTA) GNN and GT models on both homophilic and heterophilic graphs, where nearby nodes tend to have the same or different labels. Besides, we demonstrate the scalability of Polynormer on large-scale graphs that contain millions of nodes. In addition, we perform additional ablation analysis to understand the effectiveness of the local-to-global attention scheme adopted by Polynormer. Finally, we visualize Polynormer attention scores to showcase its ability of capturing critical global structures.
Experimental Setup. Appendix E shows the details of all $13$ datasets used in our experiments, which consist of $7$ homophilic and $6$ heterophilic graphs. Notably, the heterophilic datasets are from Platonov et al. (2023); Lim et al. (2021), which have addressed serious issues (e.g., train-test data leakage) of conventional popular datasets such as chameleon and squirrel. We compare Polynormer against $10$ competitive GNNs. Note that prior graph polynomial models except tGNN (Hua et al., 2022) discussed in Section 2 run out of GPU memory even on the smallest dataset in our experiments. Besides, we also evaluate $6$ GTs that have shown promising results on the node classification task. We report the performance results of baselines from their original papers or official leaderboards whenever possible, as those results are obtained by well-tuned models. For baselines whose results are not publicly available on given datasets, we tune their hyperparameters to achieve the highest possible accuracy. Detailed hyperparameter settings of baselines and Polynormer are available in Appendix H. Our hardware information is provided in Appendix F.
@@ 4.1 Performance on Homophilic and Heterophilic Graphs
Performance on homophilic graphs. We first compare Polynormer with $16$ popular baselines that include SOTA GNNs or GTs on $5$ common homophilic datasets. As shown in Table 1, Polynormer is able to outperform all baselines on $3$ out of $5$ datasets, which clearly demonstrate the efficacy of the proposed polynomial-expressive model. Moreover, incorporating the $ReLU$ function can lead to further improvements in the accuracy of Polynormer. This is because the nonlinearity imposed by $ReLU$ introduces additional higher-order monomials, which in turn enhance the quality of node representations. It is noteworthy that the accuracy gain of Polynormer over SOTA baselines is $1.65\%$ on Computer, which is a nontrivial improvement given that those baselines have been finely tuned on these well-established homophilic datasets.
Performance on heterophilic graphs. Table 2 reports the average results over $10$ runs on heterophilic graphs. Notably, the GT baselines underperform SOTA GNNs on all $5$ datasets, which raises concerns about whether those prior GT models properly utilize the expressivity brought by the self-attention module. Moreover, there is no baseline model that consistently ranks among the top $3$ models across $5$ datasets. In contrast, by integrating the attention mechanism into the polynomial-expressive model, Polynormer surpasses all baselines on $4$ out of $5$ datasets. Furthermore, Polynormer-r consistently outperforms baselines across all datasets with the accuracy improvement by a margin up to $3.55\%$ (i.e., $97.46\%$ vs. $93.91\%$ on minesweeper).
@@ 4.2 Performance on Large Graphs
We conduct experiments on large graphs by comparing Polynormer against GTs as well as some representative GNN models. Since the graphs in ogbn-products and pokec are too large to be loaded into the GPU memory for full-batch training, we leverage the random partitioning method adopted by prior GT models Wu et al. (2022; 2023) to perform mini-batch training. As shown in Table 3, Polynormer-r outperforms all baselines on these large graphs with the accuracy gain of up to $4.06\%$. Besides, we can observe that the accuracy of Polynormer drops $1.64\%$ when removing $ReLU$ activation on ogbn-arxiv. Thus, we believe the nonlinearity associated with $ReLU$ is more critical on ogbn-arxiv, which is known to be a challenging dataset (Shirzad et al., 2023). In addition, Appendix J provides the training time and memory usage of Polynormer and GT baselines. Moreover, we further evaluate Polynormer on an industrial-level graph benchmark named TpuGraphs (Phothilimthana et al., 2024), whose results are provided in Appendix L.
@@ 4.3 Ablation Analysis on Polynormer Attention Schemes
Figure 3 shows the comparison of the SOTA baseline (red bar), Polynormer without global attention (orange bar), a variant of Polynormer where the local and global attention modules are applied simultaneously to update node features per layer (blue bar), and Polynormer (green bar). We provide the detailed architecture of the Polynormer variant in Appendix G. Besides, we omit the results of Polynormer without local attention, as it completely ignores the graph structural information and thus performs poorly on the graph datasets.
@@ Appendix C Polynomial Expressivity of Prior Graph Models
Conclusion. Due to the limitations of polynomial expressivity analyzed above, previous GTs and high-order GNNs still necessitate nonlinear activation functions to attain reasonable accuracy. In contrast, our method achieves higher accuracy than these models even without activation functions, thanks to its high polynomial expressivity.
@@ H.2 Polynormer
Warm-up Epochs Local-to-Global Epochs Local Layers Global Layers Dropout
@@
Photo $200$ $1000$ $7$ $2$ $0.7$
CS $100$ $1500$ $5$ $2$ $0.3$
Physics $100$ $1500$ $5$ $4$ $0.5$
WikiCS $100$ $1000$ $7$ $2$ $0.5$
@@
questions $200$ $1500$ $5$ $3$ $0.2$
ogbn-arxiv $2000$ $500$ $7$ $2$ $0.5$
ogbn-products $1000$ $500$ $10$ $2$ $0.5$
pokec $2000$ $500$ $7$ $2$ $0.2$
Like the baseline models, we set the hidden dimension to $512$ and the learning rate to $0.001$. Additionally, we introduce a warm-up stage dedicated to training the local module. This step ensures that the node representations generated by the local module capture meaningful graph structural information before being passed to the global module. Moreover, we leverage $8$ heads for our attention modules. For the ogbn-products and pokec datasets, we use mini-batch training with a batch size of $100,000$ and $550,000$ respectively, while for the other datasets, we employ full batch training. Besides, we disable the local attention module on ogbn-arxiv and pokec datasets by replacing GAT (Veličković et al., 2017) with GCN (Kipf & Welling, 2016), which we empirically observe perform better with lower memory usage. In Table 5, we provide the critical hyperparameters of Polynormer used with each dataset.
Additional implementation details. We can rewrite Equation 7 as Equation 16:
${{\bm{X}}}={{\bm{H}}}\odot{\bm{A}}{{\bm{V}}}+\sigma(\mathbf{1}\bm{\beta^{T}})\odot{\bm{A}}{{\bm{V}}}$ (16)
To improve training stability, we adopt a LayerNorm on ${{\bm{H}}}\odot{\bm{A}}{{\bm{V}}}$. Besides, we empirically find that scaling LayerNorm$({{\bm{H}}}\odot{\bm{A}}{{\bm{V}}})$ by $1-\sigma(\mathbf{1}\bm{\beta^{T}})$ typically improves model accuracy. Thus, we implement the local attention module as shown in Equation 17, and the global attention module follows a similar way:
[…]
Appendix I provides the basic Polynormer implementation. We refer readers to our source code for the detailed model implementation and hyperparameter settings.
@@ Appendix I Model Implementation
To validate the linear complexity of Polynormer with respect to the number of nodes/edges in a graph, we generate synthetic Erdos-Renyi (ER) graphs with node count ranging from $1,000$ to $200,000$. We control the edge probability in the ER model to achieve an average node degree of around $5$ and set the node feature dimension to $100$. Figure 5 (left) clearly shows that the memory usage of Polynormer linearly increases with graph size, confirming its linear space complexity. As for the time complexity of Polynormer, we observe in Figure 5 (right) that the GPU training time remains nearly constant when increasing the graph size. We attribute it to that the GPU device (RTX A6000) used in our experiments can efficiently handle parallel computations on graphs with different sizes for full batch training. This observation is further supported by the CPU training time results, which linearly increase with graph size, as the CPU is less powerful than the GPU for parallel computation.
@@
ogbn-arxiv ogbn-products pokec Method Train/Epoch (s) Mem. (GB) Train/Epoch (s) Mem. (GB) Train/Epoch (s) Mem. (GB) GAT 0.16 10.64 0.97 9.04 1.36 9.57 GraphGPS $\underline{1.32}$ $\underline{38.91}$ $OOM$ $OOM$ $OOM$ $OOM$ NAGphormer $4.26$ 5.15 $9.64$ 7.91 $38.32$ 6.12 Exphormer 0.74 $\underline{34.04}$ $OOM$ $OOM$ $OOM$ $OOM$ NodeFormer $\underline{1.20}$ $\underline{16.30}$ $3.37$ $31.55$ $6.15$ $17.21$ DIFFormer $\underline{0.77}$ $\underline{24.51}$ 1.50 $16.24$ 3.93 $16.03$ GOAT $12.32$ 6.98 $20.11$ 9.64 $51.93$ 8.73 Polynormer 0.31 $\underline{16.09}$ 3.13 $12.93$ 2.64 $21.54$
@@
Apart from comparing Polynormer against conventional GRL models on mainstream graph datasets, we further evaluate it on TpuGraphs (Phothilimthana et al., 2024), a large-scale runtime prediction dataset on tensor computation graphs: github.com/google-research-datasets/tpu_graphs.
@@
Appendix M Discussion on the Definition of Polynomial Expressivity
As graph learning models primarily learn node representations by aggregating input features from a set of nodes, they can be viewed as a multiset pooling problem with auxiliary information about the graph topology (Baek et al., 2021; Hua et al., 2022). In the context of polynomial networks on graphs, we follow a similar notion by focusing on polynomial functions that map input node features into output node representations, with the graph topology encoded into polynomial coefficients. Notably, such polynomials have been explored by prior studies Hua et al. (2022); Chrysos et al. (2020; 2022); Wang et al. (2021), which inspire us to formally provide Definition 3.1. Moreover, assessing the ability to learn high-degree polynomials is well motivated based on the Weierstrass theorem, which guarantees any smooth function can be approximated by a polynomial Stone (1948). Thus, models with higher polynomial expressivity are more capable of learning complex functions.
While Definition 3.1 does not explicitly mention graph topology, the utilization of graph topology in different graph learning models affects the inclusion of distinct monomial terms in the learned polynomial function, thereby impacting the polynomial expressivity. It is also worth noting that Definition 3.1 can be viewed as a simplified version of the definition on polynomial expressivity in Puny et al. (2023), since we consider graph structural information as polynomial coefficients rather than indeterminates. This simplification allows us to decouple the analysis on node features from graph topology, which renders us to develop a scalable model with high polynomial expressivity.